aevans182:

The number of people living in a country is increasing each year exponentially so that the number of people 5 years ago was 4 million. The number of people in five years time is projected to be 6.25 million. What is the present population of the country?

aevans182:

I have no idea what this is. Help plz!

aevans182:

@AZ

aevans182:

@Angle

aevans182:

Can someone help me plz?

aevans182:

@Angle @AZ can someone help me plz

Angle:

uh I can try, but I gotta go teach it to myself first lol here is what I'm thinking so far... We have two points: (-5, 4) and (5, 6.25) we are looking for 1. the equation 2. the value of y for (0, y)

aevans182:

lol

aevans182:

ok so how do i do that

Angle:

I googled "how to find exponential equation given two points" and now I'm reading this https://sciencing.com/exponential-equation-two-points-8117999.html

AZ:

Can't you just use the five years ago as a starting point for the exponential equation So you initially start off with 4 million and in 10 years you'll have 6.25 million and then once you get the rate, you can just find the population in 5 years which would be the current time

aevans182:

umm idek my teacher didn't explain it very good

aevans182:

would it be 5.125 million?

Angle:

oh so like... (0, 4) and (10, 6.25) and we looking for (5, y)

AZ:

Yeah, it would make it much easier to solve

Angle:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @aevans182 would it be 5.125 million? \(\color{#0cbb34}{\text{End of Quote}}\) close but not quite I think

aevans182:

oh umm 5 million?

Angle:

\(y = ab^x\) this is the format we plan on using to try to create an equation we are going to temporarily plug in (0, 4) in order to solve for a or b

Angle:

\(4 = a~b~^0\)

Angle:

what is \(b^0\) = ?

aevans182:

1?

aevans182:

im so confused

Angle:

great! so then what is "a" equal to?

Angle:

\(4 = a \times 1\)

aevans182:

4

Angle:

ok so our formula is \(y = 4~b~^x\) so far we are going to temporarily plug in the point (10, 6.25) and try to solve for b

Katrin369:

I will give an example when Angle in done.

Katrin369:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Katrin369 I will give an example when Angle in done. \(\color{#0cbb34}{\text{End of Quote}}\) An example I use.

Katrin369:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Katrin369 \(\color{#0cbb34}{\text{Originally Posted by}}\) @Katrin369 I will give an example when Angle in done. \(\color{#0cbb34}{\text{End of Quote}}\) An example I use. \(\color{#0cbb34}{\text{End of Quote}}\) That might help you.

Angle:

\(6.25 = 4 \times b~^{10}\) what would "b" be equal to? (I would suggest using a calculator)

aevans182:

i got like a 9.3 number

aevans182:

9.31322575e-10 bytes^10

Angle:

what's it say the solutions are?

aevans182:

1.04564 and -1.04564

Angle:

cool, let's just take the positive one that's what b is equal to

Angle:

now our equation is \(y = 4(1.04564)^x\)

Angle:

plug in x = 5 solve for y that would be your answer

aevans182:

so 5

Angle:

million yes :)

aevans182:

oh ok thank you so much